Previous

Calendar

April

2017

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

26

JD85

27

JD86

28

JD87
Snow (warm, wet)

from approx 16:30-17:30 MST, power outage at MRS

29

JD88
from approx 10:11-10:30 MST, short power outage at MRS

30

JD89
Site Visit (SpB): Details , Photos (30)

Span cylinder at 1400 psi. N2 at 20 --> 1800 psi

from 14:00-14:15 MST, removed cc267882 (N2) at 20 psi started cc488594 at 1800 psi

brought cc267882 back to C-1 trailer

at 16:00 MST, replaced domes on uc REBS Q*7-1 Rnet sensor; releveled the uc CNR-2 rad sensor
Problematic thermocouples: T3 (dead), T4 (dead)
SpB Jeep

Warm, calm, partially-cloudy

31

JD90
CU Holiday (Spring Break)

1

JD91

2

JD92

3

JD93

4

JD94

5

JD95

6

JD96

7

JD97

8

JD98

9

JD99

10

JD100

11

JD101
(Full "Pink" Moon)

12

JD102

13

JD103
Site Visit (SpB, KH, HB): Details , Photos (30)

Span cylinder at 1370 psi. N2 at 1600 psi

brought ccXXXXXX from C-1 trailer to the tower (in shed)

from 13:20-14:10 MST, cleaned the LI-7200 windows and cell

from 15:00-16:00 MST, swapped the 2-micron Nupro filter at LI-6262 inlet (big increase to flow!)

Problematic thermocouples: T3 (still dead), T4 (still dead)

Kriss H and Henry B took out two problematic trees near the tower (and two fallen trees on the trail)

SpB Jeep

Warm, sunny

14

JD104

15

JD105

16

JD106

17

JD107

18

JD108
GLEES Site Visit (JF, WM, SpB): Details , Photos (40)

trip to GLEES, swpped p transducers (22.65m with 6.5m)

moved FCO44 transducer (SN 9301150) and moved Qualmetrics quad-disk port from 6.5m to 22.65m

moved FCO44 transducer (SN 950329) and moved Bedard quad-disk port from 22.65m to 6.5m

Enterprise rental

Warm w/ steady wind (GLEES)

19

JD109

20

JD110

21

JD111

22

JD112
Earth Day.

23

JD113

24

JD114

25

JD115
GLEES Site Visit (JF, WM): Details , Photos (none)

moved FCO44 transducer (SN 950329) and quad-disk port from 6.5m to 22.65m

both quad-disk ports are now co-located side-by-side at 22.65m

removed voltage dividers/reducers from each transducer (so range increased from +/-5V to +/- 15V, and clipping should be expected)

NOTE: voltage divider for SN 9301150 appeared to be damaged

Snowy, windy (GLEES)

26

JD116

27

JD117
light snow

28

JD118
in PM: cold front passage

29

JD119
snow

30

JD120

1

JD121

2

JD122
GLEES Site Visit (JF): Details , Photos (none)

both quad-disk ports are still co-located side-by-side at 22.65m

added repaired voltage divider back to transducer SN 9301150 (in grey box)

Snowy (GLEES)

3

JD123

4

JD124
Site Visit (SpB): Details , Photos (26)

Span cylinder at 1170 psi. N2 at 1300 psi

from 13:52-13:55 MST, adjusted flow of N2/Span for LI-6262

cleaned old/nasty insulation out of TGA shed

Problematic thermocouples: T3 (still dead), T4 (still dead)

Enterprise rental

Warm, sunny, calm

5

JD125

6

JD126